OpenStudy (anonymous):

(1. suppose the postition of an object moving horizontally after t seconds is given by the function s = f(t) = 2t^3 - 21 t^2 + 60 t where s is measured in feet, with s > 0 corresponding to positions right of the orgin. EXPLAIN ALL THE STEPS! A) Find the velocity and acceleration of the object. B) Determine the acceleration of the object when its velocity is zero.

OpenStudy (anonymous):

@AccessDenied @zepdrix @robtobey

zepdrix (zepdrix):

The velocity function is given by the derivative of the position function s.\[\Large\sf s'(t)=v(t)\]So our s(t) is,\[\Large\sf s(t)=2t^3-21t^2+60t\]So what is the derivative?\[\Large\sf s'(t)=?\]Apply the power rule to each term.

OpenStudy (anonymous):

Can you please show me how to apply the power rule and step by step so I can understand?

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

@ayeshaafzal221 @ChristinaWiler @dumbsearch2 @elaornelas @eliassaab @ifrimpanainte @Kira_Yamato @kirbykirby

OpenStudy (anonymous):

@nitz

OpenStudy (anonymous):

The power rule is:\[d/dx(x ^{n}) = nx ^{n-1}\]

OpenStudy (anonymous):

The another result which will be used over here is that derivative of sum or difference of two functions is equal to sum or difference of their independent derivatives wrt to the variable..

OpenStudy (anonymous):

d/dx ((f(x)+g(x)) = d/dx(f(x)) + d/dx((g(x))

OpenStudy (anonymous):

Always remember if y=f(x) ,then the derivative is dy/dx ie. derivative of dependent wrt to independent variable

OpenStudy (anonymous):

dy/dx or \[f \prime (x)\] means the same thing...derivative of y wrt x

OpenStudy (anonymous):

So,\[s \prime (t) =6t ^{2}-42t +60\]

OpenStudy (anonymous):

What would the acceleration be for B?

OpenStudy (anonymous):

acceleration is \[s \prime \prime (t) \]

OpenStudy (anonymous):

So it is s'(t) = 6t^3 - 42t + 60 that's for B?

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

@washcaps @pgpilot326 @praxer @Miamy

OpenStudy (anonymous):

@RadEn @Rvc @richyw

OpenStudy (alekos):

no, its 12t - 42

OpenStudy (anonymous):

Here is a complete solution \[ s(t)=2 t^3-21 t^2+60 t\\ s'(t)=6 t^2-42 t+60\\ s''(t)=12 t-42\\ s'(t) =0 \text { for } t=2, t=5\\ s''(2)=-18\\ s''(5)=18\\ \]